What is the system of equations that represents the problem presented below
Nan's Nursery sells bushes, trees, and perennial flowers. Each bush costs $20, each tree costs $40, and each flowers costs $2. To plant these bushes, trees, and flowers, Nan's chargers $10 per bush, $10 per tree, and $1 per flower. Also Nan will insure the items she plants for 1 year at a charge of $4 per bush, $6 per tree, and $0.20 per flower. A new Homeowner has budgeted $1000 for purchases, $400 for planting, and $150 for insuring purchases from Nan's. How many bushes, trees, and flowers should the homeowner buy to exactly use up the money budgeted?

Answer by richwmiller(17219)

20b+40t+2f=1000
10b+10t+1f=400
4b+6t+.2f=150